# HET/HPF observations of Helium in warm, hot, and ultra-hot Jupiters

### v1

@@ Appendix B A note on Ca ii IRT inspection with HET/HPF
This work focuses on the He i triplet line, but the Ca ii infrared triplet (IRT; 8500 Å, 8544 Å, and 8664 Å) lines are also accessible with the HPF spectrograph. In this section, we explore HPF’s ability to detect and measure Ca ii IRT absorption signals in planetary atmospheres. In particular, we focus on WASP-76 b because N. Casasayas-Barris et al. (2021a) already reported the detection of Ca ii IRT lines combining two transits observed by the CARMENES spectrograph.
We observed two consecutive transits of WASP-76 b with HPF: visit 1 has out-of-transit baseline the night before and during the same night, and visit 2 has out-of-transit baseline one and two nights after. We find that the transmission spectra obtained using out-of-transit data from farther in time were noisier than those obtained using data from closer to the transiting night. The most reliable results are obtained from visit 1 using only the spectra taken on the same night, which is shown in Figure 8. The TS from other out-of-transit selections are strongly affected by stellar variability. Our findings are expected since the Ca ii IRT lines are used as stellar activity tracers (M. Zechmeister et al., 2018).
The 8544 Å, and 8664 Å TS exhibit 2% features, much deeper than the 0.6% and 0.35% absorptions, respectively, reported by N. Casasayas-Barris et al. (2021a). The 8500 Å TS does not have the precision to confirm the previous detection of 0.22%, and we can only to rule out absorptions greater than 1%. If the signals at 8544 Å, and 8664 Å are real, e.g., due to planetary absorption variability, then the absorption at 8500 Å should be deeper as well. However, there is no evidence for this scenario. Furthermore, the three residual maps show vertical structure at the line positions, likely originating from line core variability. The stellar Ca ii IRT lines are usually stronger and deeper than the He triplet line resulting in lower S/N at the core. The residual maps from N. Casasayas-Barris et al. (2021a, Figure 8) show residuals from the line core as well, but the planetary trace is visible at phases far from the central transit time, due to the large velocity shift of the planet. Our visit 1 with HPF only covered a fraction of the transit close to the central time, where the planetary trace intersects the core. Therefore, any planetary signal is buried in the core variability. Similar results were obtained for HAT-P-67 b. M. Gully-Santiago et al. (2024) with HPF did not detect the Ca ii IRT absorption reported by A. Bello-Arufe et al. (2023) with CARMENES.
@@
Appendix C Reference stellar spectra and individual transmission spectra
Appendix D Stellar He triplet light curves

### v2

@@ Appendix B A note on Ca ii IRT inspection with HET/HPF
This work focuses on the He i triplet line, but the Ca ii infrared triplet (IRT; 8500 Å, 8544 Å, and 8664 Å) lines are also accessible with the HPF spectrograph. In this section, we explore HPF’s ability to detect and measure Ca ii IRT absorption signals in planetary atmospheres. In particular, we focus on WASP-76 b because Casasayas-Barris et al. (2021) already reported the detection of Ca ii IRT lines combining two transits observed by the CARMENES spectrograph.
We observed two consecutive transits of WASP-76 b with HPF: visit 1 has out-of-transit baseline the night before and during the same night, and visit 2 has out-of-transit baseline one and two nights after. We find that the transmission spectra obtained using out-of-transit data from farther in time were noisier than those obtained using data from closer to the transiting night. The most reliable results are obtained from visit 1 using only the spectra taken on the same night, which is shown in Figure 8. The TS from other out-of-transit selections are strongly affected by stellar variability. Our findings are expected since the Ca ii IRT lines are used as stellar activity tracers (Zechmeister et al., 2018).
The 8544 Å, and 8664 Å TS exhibit 2% features, much deeper than the 0.6% and 0.35% absorptions, respectively, reported by Casasayas-Barris et al. (2021). The 8500 Å TS does not have the precision to confirm the previous detection of 0.22%, and we can only to rule out absorptions greater than 1%. If the signals at 8544 Å, and 8664 Å are real, e.g., due to planetary absorption variability, then the absorption at 8500 Å should be deeper as well. However, there is no evidence for this scenario. Furthermore, the three residual maps show vertical structure at the line positions, likely originating from line core variability. The stellar Ca ii IRT lines are usually stronger and deeper than the He triplet line resulting in lower S/N at the core. The residual maps from Casasayas-Barris et al. (2021, Figure 8) show residuals from the line core as well, but the planetary trace is visible at phases far from the central transit time, due to the large velocity shift of the planet. Our visit 1 with HPF only covered a fraction of the transit close to the central time, where the planetary trace intersects the core. Therefore, any planetary signal is buried in the core variability. Similar results were obtained for HAT-P-67 b, where Gully-Santiago et al. (2024) with HPF did not detect the Ca ii IRT absorption reported by Bello-Arufe et al. (2023) with CARMENES.
@@
Appendix C Stellar He triplet light curves
Appendix D Reference stellar spectra and individual transmission spectra